OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (anonymous):

I would start by finding the 12th term, right?

OpenStudy (igreen):

Well, find the rest of the terms.

OpenStudy (igreen):

Do you see the pattern?

OpenStudy (anonymous):

It's easy if you find the common difference in the pattern, and realize it will most likely be a arithmetic progression.

OpenStudy (anonymous):

Remember for common difference, \[t_2-t_1 = t_3-t_2\] where t represents term and the subscript is term in the progression

OpenStudy (anonymous):

It's -5, right?

OpenStudy (anonymous):

That's the pattern

OpenStudy (anonymous):

Right on :), but there is a short cut, to figuring out the 12th term, or 100th, or 5000th, or millionth, you get the point.

OpenStudy (anonymous):

hehe, yeah :)

OpenStudy (igreen):

Yep, we're subtracting 5.

OpenStudy (anonymous):

So four our problem here, we will use \[t_n = a+(n-1) \times d\] a is the first term in the sequence, d is the common difference and n is the term you're looking for.

OpenStudy (anonymous):

It's pretty neat, we can derive it from your problem, but do this for me, use this and than compare it by doing it the tedious way :P

OpenStudy (anonymous):

so tn = 1 + (n-1) * -5

OpenStudy (anonymous):

oops forgot 12 tn= 1 + (12 - 1) * -5

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

so -60?

OpenStudy (anonymous):

Mhm, nope.

OpenStudy (anonymous):

Try again

OpenStudy (anonymous):

oh my goodness - 54

OpenStudy (anonymous):

got my order of operations wrong :P

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

Is that the whole answer? I feel like I am missing something.. hm

OpenStudy (anonymous):

Now you can even try by doing it without the little formula we just used

OpenStudy (anonymous):

And you'll get the same answer ;)

OpenStudy (anonymous):

That's it, your 12th term is -54

OpenStudy (anonymous):

Oh, I see it's asking you to list the first 12 sums, you can do that as well, just -5, from each term, but you also have this neat formula now.

OpenStudy (anonymous):

So if you realized, the main thing here was noticing the common difference, and it's a breeze.

OpenStudy (anonymous):

@iambatman